# VL-LN Bench: Towards Long-horizon Goal-oriented Navigation with Active Dialogs

### v3

@@ I Introduction
A practical navigation agent must handle vague tasks by both planning effectively and resolving ambiguities. Active dialog offers a natural solution, allowing the agent to clarify underspecified instructions and obtain cues for efficient navigation. To study this ability, we propose Interactive Instance Goal Navigation (IIGN; as shown in Fig. 1), which extends Instance-level Object Navigation (ION, named as IGN in our paper) [ION]. In IIGN, the agent receives a basic ObjectNav [objectnav] instruction (e.g.,“Search for the <category>”), which is insufficient to uniquely identify the target instance, and must consult an oracle via dialog. Since IIGN inherently poses exploration and disambiguation challenges, it provides an ideal testbed for examining the role of dialog in navigation.
To effectively investigate and address the proposed IIGN task, this paper aims to build an Agent–Oracle interaction benchmark for dialog-enabled navigation agent training and evaluation. Prior work [thomason2020vision] collected human–human dialogs in navigation to evaluate whether agents can understand and follow instructions. However, such efforts do not assess an agent’s ability to proactively ask targeted questions. More recent approaches attempt to enable agents to generate questions, but they either limit the task to small-scale, room-level settings [gao2022dialfred] or focus narrowly on instance descriptions with limited support for long-horizon exploration [taioli2024collaborative]. Moreover, existing methods generally lack large-scale training datasets, which constrains the development of agents capable of both active exploration and informative questioning.
In contrast to prior work, we propose the Vision Language–Language Navigation (VL-LN) benchmark, which provides both a large-scale, automatically generated training dataset and a comprehensive evaluation protocol for developing and assessing dialog-enabled navigation agents in long-horizon settings. The automated data collecting pipeline comprises three steps: (1) aggregating room-level region and instance attributes from MMScan [mmscan] into unified house-level annotations; (2) pairing target instances with feasible initial positions to instantiate episodes; and (3) collecting dialog-rich trajectories using a frontier-based navigator and a scripted oracle that answers questions to support exploration and disambiguation. Through this process, we curate the first large-scale training dataset of $\sim 41$K dialog-augmented trajectories for IIGN. The VL-LN benchmark further incorporates an evaluation protocol with the scripted oracle, enabling the assessment of agents’ dialog generation capabilities without requiring human intervention. Together, the dataset and evaluation protocol establish a unified benchmark for training and evaluating agents’ abilities in both language-based querying and navigation (LN) across IIGN and IGN.
@@ II-A Goal-oriented navigation
Goal-oriented navigation requires an agent to find a specified goal in an unknown environment [ieong2025multimodal]. In the text-guided setting, it is commonly divided into Object-goal Navigation (ObjectNav) [anderson2018evaluation, objectnav] and Instance-level Object Navigation (ION) [ION]. Most prior work focuses on ObjectNav, where the instruction is unambiguous because any instance of the target category suffices. These approaches fall into two groups: training-based [campari2020exploiting, yadav2023offline, ye2021auxiliary, xie2023implicit, yin2025unigoal] and zero-shot [yokoyama2024vlfm]. Training-based methods typically use Imitation Learning [zare2024survey] or Reinforcement Learning [kaelbling1996reinforcement] to learn the observation-to-action mapping [ye2021auxiliary, xie2023implicit, yin2025unigoal], and developing vision–language aligned embeddings [campari2020exploiting, yadav2023offline] either in a single end-to-end model [yadav2023offline, ye2021auxiliary, xie2023implicit] or a modular pipeline [chen2023not, cai2024mo]. Zero-shot methods instead combine classical frontier-based exploration [yamauchi1997frontier] with the priors provided by large language [yu2023l3mvn, cai2024bridging] or vision–language models [majumdar2022zson, yokoyama2024vlfm] to score frontiers and guide exploration without task-specific training.
IGN takes a step further by requiring the agent to locate a specific instance rather than any category member, which reflects another facet of real-world needs, where users often care about particular or personalized targets. [ION, barsellotti2024personalized]. Recent instance goal navigation benchmarks [ION, barsellotti2024personalized] show that disambiguation is a central challenge, and that learned exploration policies often struggle to backtrack and recover after moving in the wrong direction. To address this problem, IGN [ION] builds an instance-level graph whose nodes encode color, material, and location features, and PSL [sun2024prioritized] proposes Prioritized Semantic Learning for IGN, which enhances semantic understanding to better identify the target. Despite these advances, real environments often contain many look-alike instances, and textual descriptions alone frequently lack the specificity needed to differentiate them [taioli2024collaborative]. Unlike prior work that treats the agent as a passive recipient of instance-level information, we enable the agent to proactively propose dialogs to refine the task specification and obtain targeted guidance, thus improving exploration efficiency.
@@ II-B Interactive Embodied Robotics
Methods Early work measures uncertainty to decide whether the agent should act or ask for help; upon receiving a query, the human or a simulated oracle returns the next best action or a shortest path to the target [24, 27]. Another line of work returns images of candidate objects and asks the human to confirm the target [29]. These designs restrict interaction to a single mode, either asking for the next action or confirming the target. With the emergence of interactive platforms [12] and datasets [thomason2020vision], agents can ask free-form, natural-language questions. RMM [26] simulates oracle answers and estimates their effect on progress to learn which questions to ask. KNOWNO [41] uses conformal prediction to decide when to seek help from a language-model planner. More recently, AIUTA [taioli2024collaborative] allows an agent to identify the target instance through open-ended, template-free dialog. Despite these advances, existing methods underemphasize efficient exploration in instance-goal navigation; guidance is typically limited to target descriptions [29, taioli2024collaborative] or a few rounds of local actions [24, 27, 26, 41]. Our approach supports both disambiguation questions and exploration-path questions, enabling guidance of long-term navigation.
Benchmarks We focus on natural-language interactive datasets. CVDN [thomason2020vision] collected human–human dialogs but lacks an oracle, making it unsuitable for online evaluation. DialFRED [gao2022dialfred] and CoIN [taioli2024collaborative] introduce oracles for online evaluation: DialFRED’s oracle provides best actions in small, single-room scenes, while CoIN’s oracle only describes the target instance. As a result, these settings make it difficult to evaluate exploration and querying in the more practical house-scale environments. In addition, existing training datasets [thomason2020vision, gao2022dialfred, taioli2024collaborative], are not large enough to support learning agents that both explore actively and converse effectively. We introduce VL-LN Bench, a large-scale dialog-augmented training set with a house-level oracle that answers both exploration and disambiguation queries, enabling efficient policy training and comprehensive evaluation. A detailed comparison of existing benchmarks is shown in Table I.
@@ III-A Task Definition
The Interactive Instance Goal Navigation (IIGN) challenges a dialog-enabled embodied agent to locate a specific instance in an unfamiliar environment. It involves two active roles: an agent and an oracle. For each episode, the agent is randomly placed in an unknown environment [szot2021habitat] and given an ambiguous instruction providing only the target category (e.g., “Search for the chair”). At each step, the agent receives a visual observation $o_{t}$ and odometry $l_{t}$ from the environment, and chooses to either move from $a_{t}\in\mathcal{A}$, where
@@
DialFRED [gao2022dialfred] - 53,000 room - Human ✗ / ✓ ✓(dir.)
@@ III-C1 Scenes Metadata Processing
We process the MP3D scene meta-annotations using the hierarchical labels from MMScan [mmscan], which provide fine-grained descriptions at both object and region levels. Object-level annotations include spatial properties (geometry, pose) and attributes (category, appearance, material, state, functional use), while room-level annotations include each region’s function (e.g., dining room, study, bathroom) and the objects contained in that region. Based on these annotations, we reconstruct house-level meta-annotations into an instance dictionary and a region dictionary for each scene. The key difference from MMScan is that we merge room-level annotations into a single house-level index that covers all rooms in the scene. Moreover, we build a spatial-relation graph using Sr3D [Sr3d]. Each node corresponds to an instance, and edges connect an instance to nearby instances within $1\,\mathrm{m}$. Together with the house-level dictionaries, this graph provides robust relational cues that help disambiguate instances even when they appear visually identical.
@@ III-C2 Episodes Generation
Each episode is defined by three core elements: an initial agent pose, a navigation instruction, and a set of target-instance viewpoints. The initial agent poses include starting poses from R2R-CE (72 scenes) [vln-ce] and our manually annotated 18 scenes. For the self-annotated scenes, we randomly sample navigable points, snap the agent to the selected location, and accept it as an initial pose only after verifying that the location is valid (on the navmesh, collision-free, and within scene bounds). For every instance, we provide two instruction variants: a partial, category-only instruction mentioned above and a full description that uniquely identifies the target among all instances in the scene (e.g., “Locate the deep grey chair with black backrest, standing upright on a wooden floor near a computer and a tv in the bedroom.”). To create the full description, we leverage the house-level dictionaries and the spatial-relation graph to select discriminative attributes and relations for the instance, and then prompt GPT-4o [gpt4o] to generate a natural-language instruction. The full instruction supports a non-interactive instance goal navigation setting (i.e. IGN), while the partial instruction is used in IIGN. We generate a set of viewpoints for each instance, which serve as expected “Stop” locations. Based on empirically reasonable viewing distances, we expand each instance’s 3D bounding box by $0.6\,\mathrm{m}$ in all directions and mark all navigable points within the expanded region as instance viewpoints. The episode is considered to be successful if the agent stops within 0.25 $m$ of any viewpoint of the instance. Combining an initial agent pose with the instance’s instruction and its viewpoints defines an episode.
@@ III-D Statistics
Our dataset covers 112 object categories that can serve as navigation targets (excluding structural elements such as walls and ceilings), comprising 20476 object instances across 90 annotated MP3D scenes. For each scene, we sample a set of navigable start poses (mean 42 per scene). Pairing these starts with the annotated instances yields 333,319 episodes. Following the VLN-CE partitioning [vln-ce], we allocate 61/15/14 scenes to train/val/test, resulting in 246,433 training episodes, 86,386 validation episodes, and 500 test episodes. Shown in Tab. I, we collect 41,891 trajectories spanning all scenes and categories of IIGN, each coupling navigation with question–answer interactions. Fig. 3 summarizes the trajectory statistics. In addition, we collect 5,087 IGN and 23,774 ObjectNav trajectories generated by an FBE agent for the experiment.
@@ IV-A Experimental Setup
Baselines We evaluate five baselines: two zero-shot and three training-based methods. The zero-shot baselines are (i) a greedy frontier-based exploration (FBE) agent that repeatedly selects the nearest frontier and uses an open-vocabulary detector built on Grounded SAM 2 [GroundingSam2] to detect the target instance, and (ii) VLFM [yokoyama2024vlfm], used as the released version. The learning-based baselines (VLLN-O, VLLN-I, VLLN-D) are initialized from Qwen2.5-VL-7B-Instruct and trained following the InternVLA-N1 [internvla-n1] procedure but with different data mixtures. All three include the VLN data from InternVLA-N1. VLLN-O additionally uses our ObjectNav data (23,774 trajectories), VLLN-I and VLLN-D further incorporate the filtered IGN data (11,661 trajectories): VLLN-I trained without dialog, whereas VLLN-D trained with dialogs.
@@ IV-B Result
Dialog Helps. Across both IIGN and IGN, VLLN-D achieves the best results, demonstrating the value of proactive querying while leaving room for further improvement. Dialog in the ambiguity setting (IIGN) is more effective than in the disambiguation setting (IGN) (MSP: 2.76 vs. 2.16). To better analyze the role of dialog during instance goal navigation, we categorize failures into three types: Exploration Fail, Detection Fail, and Stop Fail. Exploration Fail occurs when the agent keeps acting until the maximum number of steps without ever entering the oracle-success radius ($OS=0$). Unlike prior goal-oriented navigation settings, we further split Detection Fail into Wrong Detection and Ambiguity: the former happens when the agent reaches the correct region ($OS=1$) but fails to recognize the target instance, while the latter occurs when it stops at an instance of the same category as the target. Stop Fail denotes cases where the agent stops near the target (<1m<1m) but remains beyond the stop threshold (>0.25m>0.25m). As shown in Tab. III, when the agent is allowed to query an oracle, exploration failures decrease substantially: from 89 to 71 in IIGN and from 84 to 46 in IGN. This demonstrates that the agent can leverage information obtained through dialog to improve exploration efficiency. Notably, as shown in the supplementary video, the agent is able to ground natural-language short-route guidance into executable actions. We attribute it to co-training with VLN data, which, although different from goal-oriented navigation, teaches the model to follow natural-language route descriptions.
@@
Moreover, dialog helps resolve ambiguity. In both IIGN and IGN settings, ambiguity-related failures decrease once the dialog is enabled. In IGN, the reduction from 143 to 124 primarily stems from disambiguation questions that guide the agent away from incorrect instances, as the attributes necessary to verify the target instance are already available, rendering additional attribute-related queries less informative. For IIGN, both attribute and disambiguation questions contribute to eliminating task ambiguity, make the ambiguity failures drop from 143 to 124.
@@
Key Challenges and Directions. To better understand IIGN’s challenges, we additionally evaluate four interaction settings on 100 episodes randomly sampled from the test set: Human-Human, Human-Oracle, Agent-Human, and Agent-Oracle. Results are reported in Tab. IV. First, we argue that image–attribute alignment is the primary bottleneck in both the IGN and IIGN tasks. Even with full instructions and dialog capability, the success rate remains 25% (shown in Tab. II), and 73% of failures (shown in Tab. III) are due to detection errors. A promising direction is to train with hard negatives such as same-category objects with different attributes, which promotes instance-level discrimination and stronger attribute grounding. Zero-shot methods face the same challenge. Their performance hinges on detectors that can align detailed instance descriptions with visual evidence.
@@
Navigator–Oracle SR $\uparrow$ SPL $\uparrow$ OS $\uparrow$ NE $\downarrow$ Avg. turns
@@
Human–Oracle $91$ $49.88$ $94$ $0.69$ $9.72$
Agent–Human $16$ $12.63$ $55$ $7.02$ $1.54$
Agent–Oracle $17$ $12.05$ $55$ $8.04$ $1.66$
Figure 4: Failure cases. Green curves denote the geodesic shortest paths; blue curves are the navigator’s exploration trajectories; red shaded regions indicate the success zone around the target. (a) Referential ambiguity: within the same view, the navigator and the Oracle refer to different instances, causing the navigator to stop at a wrong instance. (b) Partial observability: the navigator only observes a single candidate in the room and stops without disambiguating. (c) Exploration failure: despite continued interaction, the human navigator never finds the target.
Secondly, the agent’s questioning ability remains limited. Reliably disambiguating a target instance from same-category distractors through dialog is still difficult. VLLN-D performs worse on IIGN (20.2%) than VLLN-I on IGN (22.4%), indicating that dialog offers less guidance than a full instruction. We also find that a gap persists between agent-driven and human-driven proactive interaction. Human-Human reaches 93% SR with only about two questions on average, which suggests that a small number of well-chosen queries is sufficient to complete the task. Achieving such behavior requires the agent to exploit its observation history to select maximally informative questions that either guide exploration or shrink the candidate set, placing stronger demands on grounding, planning, and reasoning.

### v4

@@ I Introduction
A practical navigation agent must handle vague tasks by both planning effectively and resolving ambiguities. Active dialog offers a natural solution, allowing the agent to clarify underspecified instructions and obtain cues for efficient navigation. To study this ability, we propose Interactive Instance Goal Navigation (IIGN; as shown in Fig. 1), which extends Instance-level Object Navigation (ION, named as IGN in our paper) [1]. In IIGN, the agent receives a basic ObjectNav [2] instruction (e.g.,“Search for the <category>”), which is insufficient to uniquely identify the target instance, and must consult an oracle via dialog. Since IIGN inherently poses exploration and disambiguation challenges, it provides an ideal testbed for examining the role of dialog in navigation.
To effectively investigate and address the proposed IIGN task, this paper aims to build an Agent–Oracle interaction benchmark for dialog-enabled navigation agent training and evaluation. Prior work [3] collected human–human dialogs in navigation to evaluate whether agents can understand and follow instructions. However, such efforts do not assess an agent’s ability to proactively ask targeted questions. More recent approaches attempt to enable agents to generate questions, but they either limit the task to small-scale, room-level settings [4] or focus narrowly on instance descriptions with limited support for long-horizon exploration [5]. Moreover, existing methods generally lack large-scale training datasets, which constrains the development of agents capable of both active exploration and informative questioning.
In contrast to prior work, we propose the Vision Language–Language Navigation (VL-LN) benchmark, which provides both a large-scale, automatically generated training dataset and a comprehensive evaluation protocol for developing and assessing dialog-enabled navigation agents in long-horizon settings. The automated data collecting pipeline comprises three steps: (1) aggregating room-level region and instance attributes from MMScan [6] into unified house-level annotations; (2) pairing target instances with feasible initial positions to instantiate episodes; and (3) collecting dialog-rich trajectories using a frontier-based navigator and a scripted oracle that answers questions to support exploration and disambiguation. Through this process, we curate the first large-scale training dataset of $\sim 41$K dialog-augmented trajectories for IIGN. The VL-LN benchmark further incorporates an evaluation protocol with the scripted oracle, enabling the assessment of agents’ dialog generation capabilities without requiring human intervention. Together, the dataset and evaluation protocol establish a unified benchmark for training and evaluating agents’ abilities in both language-based querying and navigation (LN) across IIGN and IGN.
@@ II-A Goal-oriented navigation
Goal-oriented navigation requires an agent to find a specified goal in an unknown environment [7]. In the text-guided setting, it is commonly divided into Object-goal Navigation (ObjectNav) [8, 2] and Instance-level Object Navigation (ION) [1]. Most prior work focuses on ObjectNav, where the instruction is unambiguous because any instance of the target category suffices. These approaches fall into two groups: training-based [9, 10, 11, 12, 13] and zero-shot [14]. Training-based methods typically use Imitation Learning [15] or Reinforcement Learning [16] to learn the observation-to-action mapping [11, 12, 13], and developing vision–language aligned embeddings [9, 10] either in a single end-to-end model [10, 11, 12] or a modular pipeline [17, 18]. Zero-shot methods instead combine classical frontier-based exploration [19] with the priors provided by large language [20, 21] or vision–language models [22, 14] to score frontiers and guide exploration without task-specific training.
IGN takes a step further by requiring the agent to locate a specific instance rather than any category member, which reflects another facet of real-world needs, where users often care about particular or personalized targets. [1, 23]. Recent instance goal navigation benchmarks [1, 23] show that disambiguation is a central challenge, and that learned exploration policies often struggle to backtrack and recover after moving in the wrong direction. To address this problem, IGN [1] builds an instance-level graph whose nodes encode color, material, and location features, and PSL [24] proposes Prioritized Semantic Learning for IGN, which enhances semantic understanding to better identify the target. Despite these advances, real environments often contain many look-alike instances, and textual descriptions alone frequently lack the specificity needed to differentiate them [5]. Unlike prior work that treats the agent as a passive recipient of instance-level information, we enable the agent to proactively propose dialogs to refine the task specification and obtain targeted guidance, thus improving exploration efficiency.
@@ II-B Interactive Embodied Robotics
Methods Early work measures uncertainty to decide whether the agent should act or ask for help; upon receiving a query, the human or a simulated oracle returns the next best action or a shortest path to the target [25, 26]. Another line of work returns images of candidate objects and asks the human to confirm the target [27]. These designs restrict interaction to a single mode, either asking for the next action or confirming the target. With the emergence of interactive platforms [28] and datasets [3], agents can ask free-form, natural-language questions. RMM [29] simulates oracle answers and estimates their effect on progress to learn which questions to ask. KNOWNO [30] uses conformal prediction to decide when to seek help from a language-model planner. More recently, AIUTA [5] allows an agent to identify the target instance through open-ended, template-free dialog. Despite these advances, existing methods underemphasize efficient exploration in instance-goal navigation; guidance is typically limited to target descriptions [27, 5] or a few rounds of local actions [25, 26, 29, 30]. Our approach supports both disambiguation questions and exploration-path questions, enabling guidance of long-term navigation.
Benchmarks We focus on natural-language interactive datasets. CVDN [3] collected human–human dialogs but lacks an oracle, making it unsuitable for online evaluation. DialFRED [4] and CoIN [5] introduce oracles for online evaluation: DialFRED’s oracle provides best actions in small, single-room scenes, while CoIN’s oracle only describes the target instance. As a result, these settings make it difficult to evaluate exploration and querying in the more practical house-scale environments. In addition, existing training datasets [3, 4, 5], are not large enough to support learning agents that both explore actively and converse effectively. We introduce VL-LN Bench, a large-scale dialog-augmented training set with a house-level oracle that answers both exploration and disambiguation queries, enabling efficient policy training and comprehensive evaluation. A detailed comparison of existing benchmarks is shown in Table I.
@@ III-A Task Definition
The Interactive Instance Goal Navigation (IIGN) challenges a dialog-enabled embodied agent to locate a specific instance in an unfamiliar environment. It involves two active roles: an agent and an oracle. For each episode, the agent is randomly placed in an unknown environment [31] and given an ambiguous instruction providing only the target category (e.g., “Search for the chair”). At each step, the agent receives a visual observation $o_{t}$ and odometry $l_{t}$ from the environment, and chooses to either move from $a_{t}\in\mathcal{A}$, where
@@
DialFRED [4] - 53,000 room - Human ✗ / ✓ ✓(dir.)
@@ III-C1 Scenes Metadata Processing
We process the MP3D scene meta-annotations using the hierarchical labels from MMScan [6], which provide fine-grained descriptions at both object and region levels. Object-level annotations include spatial properties (geometry, pose) and attributes (category, appearance, material, state, functional use), while room-level annotations include each region’s function (e.g., dining room, study, bathroom) and the objects contained in that region. Based on these annotations, we reconstruct house-level meta-annotations into an instance dictionary and a region dictionary for each scene. The key difference from MMScan is that we merge room-level annotations into a single house-level index that covers all rooms in the scene. Moreover, we build a spatial-relation graph using Sr3D [32]. Each node corresponds to an instance, and edges connect an instance to nearby instances within $1\,\mathrm{m}$. Together with the house-level dictionaries, this graph provides robust relational cues that help disambiguate instances even when they appear visually identical.
@@ III-C2 Episodes Generation
Each episode is defined by three core elements: an initial agent pose, a navigation instruction, and a set of target-instance viewpoints. The initial agent poses include starting poses from R2R-CE (72 scenes) [33] and our manually annotated 18 scenes. For the self-annotated scenes, we randomly sample navigable points, snap the agent to the selected location, and accept it as an initial pose only after verifying that the location is valid (on the navmesh, collision-free, and within scene bounds). For every instance, we provide two instruction variants: a partial, category-only instruction mentioned above and a full description that uniquely identifies the target among all instances in the scene (e.g., “Locate the deep grey chair with black backrest, standing upright on a wooden floor near a computer and a tv in the bedroom.”). To create the full description, we leverage the house-level dictionaries and the spatial-relation graph to select discriminative attributes and relations for the instance, and then prompt GPT-4o [34] to generate a natural-language instruction. The full instruction supports a non-interactive instance goal navigation setting (i.e. IGN), while the partial instruction is used in IIGN. We generate a set of viewpoints for each instance, which serve as expected “Stop” locations. Based on empirically reasonable viewing distances, we expand each instance’s 3D bounding box by $0.6\,\mathrm{m}$ in all directions and mark all navigable points within the expanded region as instance viewpoints. The episode is considered to be successful if the agent stops within 0.25 $m$ of any viewpoint of the instance. Combining an initial agent pose with the instance’s instruction and its viewpoints defines an episode.
@@ III-D Statistics
Our dataset covers 112 object categories that can serve as navigation targets (excluding structural elements such as walls and ceilings), comprising 20476 object instances across 90 annotated MP3D scenes. For each scene, we sample a set of navigable start poses (mean 42 per scene). Pairing these starts with the annotated instances yields 333,319 episodes. Following the VLN-CE partitioning [33], we allocate 61/15/14 scenes to train/val/test, resulting in 246,433 training episodes, 86,386 validation episodes, and 500 test episodes. Shown in Tab. I, we collect 41,891 trajectories spanning all scenes and categories of IIGN, each coupling navigation with question–answer interactions. Fig. 3 summarizes the trajectory statistics. In addition, we collect 5,087 IGN and 23,774 ObjectNav trajectories generated by an FBE agent for the experiment.
@@ IV-A Experimental Setup
Baselines We evaluate five baselines: two zero-shot and three training-based methods. The zero-shot baselines are (i) a greedy frontier-based exploration (FBE) agent that repeatedly selects the nearest frontier and uses an open-vocabulary detector built on Grounded SAM 2 [35] to detect the target instance, and (ii) VLFM [14], used as the released version. The learning-based baselines (VLLN-O, VLLN-I, VLLN-D) are initialized from Qwen2.5-VL-7B-Instruct and trained following the InternVLA-N1 [36] procedure but with different data mixtures. All three include the VLN data from InternVLA-N1. VLLN-O additionally uses our ObjectNav data (23,774 trajectories), VLLN-I and VLLN-D further incorporate the filtered IGN data (11,661 trajectories): VLLN-I trained without dialog, whereas VLLN-D trained with dialogs.
@@ IV-B Result
Dialog Helps. Across both IIGN and IGN, VLLN-D achieves the best results, demonstrating the value of proactive querying while leaving room for further improvement. Dialog in the ambiguity setting (IIGN) is more effective than in the disambiguation setting (IGN) (MSP: 2.76 vs. 2.16). To better analyze the role of dialog during instance goal navigation, we categorize failures into three types: Exploration Fail, Detection Fail, and Stop Fail. Exploration Fail occurs when the agent keeps acting until the maximum number of steps without ever entering the oracle-success radius ($OS=0$). Unlike prior goal-oriented navigation settings, we further split Detection Fail into Wrong Detection and Ambiguity: the former happens when the agent reaches the correct region ($OS=1$) but fails to recognize the target instance, while the latter occurs when it stops at an instance of the same category as the target. Stop Fail denotes cases where the agent stops near the target (<1m<1m) but remains beyond the stop threshold (>0.25m>0.25m). As shown in Tab. III, when the agent is allowed to query an NPC, exploration failures decrease substantially: from 89 to 71 in IIGN and from 84 to 46 in IGN. This demonstrates that the agent can leverage information obtained through dialog to improve exploration efficiency. Notably, as shown in the supplementary video, the agent is able to ground natural-language short-route guidance into executable actions. We attribute it to co-training with VLN data, which, although different from goal-oriented navigation, teaches the model to follow natural-language route descriptions.
@@
Moreover, dialog helps resolve ambiguity. In both IIGN and IGN settings, ambiguity-related failures decrease once the dialog is enabled. In IGN, the reduction from 143 to 124 primarily stems from disambiguation questions that guide the agent away from incorrect instances, as the attributes necessary to verify the target instance are already available, rendering additional attribute-related queries less informative. For IIGN, both attribute and disambiguation questions contribute to eliminating task ambiguity, make the ambiguity failures drop from 159 to 145 .
@@
Key Challenges and Directions. To better understand IIGN’s challenges, we additionally evaluate four interaction settings on 100 episodes randomly sampled from the test set: Human - Human, Human - GPT, VLLN-D - Human, and VLLN-D - GPT. Results are reported in Tab. IV. First, we argue that image–attribute alignment is the primary bottleneck in both the IGN and IIGN tasks. Even with full instructions and dialog capability, the success rate remains 25% (shown in Tab. II), and 73% of failures (shown in Tab. III) are due to detection errors. A promising direction is to train with hard negatives such as same-category objects with different attributes, which promotes instance-level discrimination and stronger attribute grounding. Zero-shot methods face the same challenge. Their performance hinges on detectors that can align detailed instance descriptions with visual evidence.
@@
Navigator–NPC SR $\uparrow$ SPL $\uparrow$ OS $\uparrow$ NE $\downarrow$ Avg. turns
@@
Human – GPT $91$ $49.88$ $94$ $0.69$ $9.72$
VLLN-D – Human $16$ $12.63$ $55$ $7.02$ $1.54$
VLLN-D – GPT $17$ $12.05$ $55$ $8.04$ $1.66$
Figure 4: Failure cases. Green curves denote the geodesic shortest paths; blue curves are the navigator’s exploration trajectories; red shaded regions indicate the success zone around the target. (a) Referential ambiguity: within the same view, the navigator and the NPC refer to different instances, causing the navigator to stop at a wrong instance. (b) Partial observability: the navigator only observes a single candidate in the room and stops without disambiguating. (c) Exploration failure: despite continued interaction, the human navigator never finds the target.
Secondly, the agent’s questioning ability remains limited. Reliably disambiguating a target instance from same-category distractors through dialog is still difficult. VLLN-D performs worse on IIGN (20.2%) than VLLN-D* on IGN (21.8%), indicating that dialog offers less guidance than a full instruction. We also find that a gap persists between agent-driven and human-driven proactive interaction. Human-Human reaches 93% SR with only about two questions on average, which suggests that a small number of well-chosen queries is sufficient to complete the task. Achieving such behavior requires the agent to exploit its observation history to select maximally informative questions that either guide exploration or shrink the candidate set, placing stronger demands on grounding, planning, and reasoning.